OpenStudy (ammarah):

Graph the function. Label the vertex and axis of symmetry. y= -2(x-4)(x-2)

OpenStudy (ammarah):

@Easyaspi314

OpenStudy (anonymous):

we do know that x = 2 and x = 4 will be the x-intercepts because when you solve -2(x-2)(x-4) = 0, you get x = 2 and x = 4.

OpenStudy (anonymous):

Now, y = -2(x-4)(x-2) = -2(x^2 - 6x + 8) = -2x^2 + 12x - 16 To find the vertex, we will compute -b/2a, that will give us the x-coordinate of the vertex. a = -2 and b = 12, so -b/2a = -12/-4 = 3 So x = 3 is the x-coordinate. To find the y-coordinate we plug in x = 3 into the equation, and we get y = 2 So the VERTEX is at (3,2) And, x = 3 will be the line of symmetry. (Note, to find line of symmetry for the parabola, it is x = -b/2a). So now we can graph this parabola. I will try graphing it.

OpenStudy (anonymous):



OpenStudy (anonymous):

Notice that (3,2) is the vertex..notice the x-intercepts of x = 2 and x = 4.

OpenStudy (anonymous):

Draw the vertical line x = 3, you will see it is a line of symmetyry (it is a line that cuts the parabola in half). Does this all make sense? Understand???

OpenStudy (anonymous):

???????????

OpenStudy (ammarah):

wait so the other points r 4, 2?